OpenStudy (anonymous):

what is the solution to |3x+10|=10 A.x=20/3 B.x=20/3,0 C.x=-20/3,0 D.x=0

hartnn (hartnn):

when |a| = b, we have a = b or a = -b so for |3x+10| = 10, we will have 3x +10 = 10 , 3x +10 = -10 solve these 2 equations to get 2 values of x

hartnn (hartnn):

can u try?

OpenStudy (anonymous):

for 3x+10=10 i got 0 and for 3x+10=-10 i got -20/3 so would that mean that the answer is C??? @hartnn

hartnn (hartnn):

correct! :)

OpenStudy (anonymous):

thank you so much

hartnn (hartnn):

welcome ^_^